OpenStudy (anonymous):

write and equations of the line that passes through the given point and is parallel to the given line. 4.(6,8) y=-5/2x+10

OpenStudy (anonymous):

8=-5/2*6+b +15 23=b y=-5/2x+23

OpenStudy (anonymous):

okay thanks what about (-1,2), y=5x+4

OpenStudy (anonymous):

Well, you would plug in what you know. -1 is x and 2 is y. You would leave b alone. If you want the same slope, then the m stays the same so: 2=5(-1)+b +5 7=b y=5x+7

OpenStudy (anonymous):

wait in the first one how did you go from 23=6+b to 23=b?